Child pages
  • 201107 - Public report
Skip to end of metadata
Go to start of metadata

You are viewing an old version of this page. View the current version.

Compare with Current View Page History

« Previous Version 4 Next »

You can manually edit this page, add pictures and so on. The content below is marked as an excerpt (see WIKI markup) and this is how the automated script will know which part of the page to update. We can also show all the tables in one page if needed.

Unscheduled:

  • 21.06.2011 14:44 UTC - A router rebooted spontaneously and the issue was raised with JTAC. NO RFO was found by JTAC since the problems never reoccurred. This ticket will be closed.
  • 13.07.2011 12:04 UTC - There was an outage on the link connecting KTHNOC to SUNET. The cause has not been found.
  • 04.07.2011 11:15 UTC - There were recurrent outages in the South Red network. The problem was fixed by changing the fibre between two amplifier sites.
  • 19.07.2011 13:38 UTC - The link between KTH and NDGF through SUNET started to flap. The flaps stopped after a patch cable was pulled out and plugged in.
  • 11.07.2011 18:53 UTC - The signal on PDCs link to DEISA was degraded and had a high error count. An intervention was held to replace a transmission card which solved the problem.
  • 14.07.2011 13:29 UTC - There was an outage on PDCs link to DEISA caused by a faulty transmission board.
  • 05.07.2011 08:38 UTC - A brief outage (2 min) was detected between Skövde and SLU Skara. We could not find any reason for the outage.
  • 02.07.2011 09:44 UTC - A brief outage between SUNET and SFHM was detected. No RFO was found.

Scheduled:

  • 04.07.2011 08:00 UTC - A planned maintenance to upgrade the Adobe Connect system was performed successfully.
  • 07.07.2011 15:14 UTC - Planned maintenance was performed on the South Red network in order to troubleshoot link flaps.
  • 11.07.2011 15:33 UTC - Fiber work was performed to troubleshoot the link outages on the South Red network.
  • 19.07.2011 13:34 UTC - A transmission card was replaced in AG_421_KTH in order to fix the performance issues between KTH and DEISA.
  • 04.07.2011 02:57 UTC - Our supplier performed reconnections of fibers in Borlänge.
  • 18.07.2011 15:06 UTC - A planned maintenance was performed trying to fix the network issues in the south red network.
  • 21.07.2011 15:30 UTC - A planned emergency maintenance was performed in order to trying fix the issues in the south red network.

Scope

Unscheduled

Scheduled

Total

Hardware

1 pcs 00:03

0 pcs

1 pcs 00:03

Link

5 pcs 06:18

6 pcs 01:10

11 pcs 07:28

None

3 pcs 1d 04:57

0 pcs

3 pcs 1d 04:57

Routing

0 pcs

0 pcs

0 pcs

Software

0 pcs

1 pcs 2d 02:00

1 pcs 2d 02:00

  • No labels